# Learning Coherent Matrixized Representation in Latent Space for Volumetric 4D Generation

### v2

@@ 1 Introduction
Recent works [13, 35, 5, 3, 40] combine the controllability of dynamic 3D objects with the expressivity of emerging diffusion models. 4D representations are proposed to construct an underlying representation of dynamic 3D objects, allowing the diffusion model to learn the distribution of 3D synthesis with an additional temporal dimension. In general, 4D representations can be categorized into prior-based and free-form methods depending on the 3D representation of the output shape. Prior-based methods are mostly derived from shape parameters combined with a series of skeletal pose parameters to model dynamic sequences, e.g., 4D human motion generatioin [17, 53]. Although they produce plausible results, the generated skeleton human motions assuming a human model prior such as SMPL [26], lacks motion diversity and only supports limited time span. Free-form methods leveraging dynamic Neural Radiance Fields (NeRF) combine the benefits of video and 3D generative models [18, 40, 6, 51], and require heavy computations to generate a 4D NeRF. Moreover, their generated motions are not well-controllable. The 4D Gaussian splatting-based free-form method [35] can produce high-quality renderings, but artifacts often appear when the view changes. This is due to the lack of geometric priors in 4D Gaussian splatting-based free-form approachs. In addition, free-form methods often produce low-quality rendering results during arbitrary navigation. This issue arises because the henerated 4D content lacks the essential volumetric properties needed for accurate representation. To this end, we pose a question, how can we efficiently represent a 4D sequence to generate with high-quality geometry, sequence consistency, and alignment with input guidance, while maintaining diversity and allowing for free navigation and rendering?
@@
To address these challenges, we propose several techniques to achieve volumetric 4D sequence generation under given conditions. We first design a coherent 3D shape and color modeling method to encode the shape and color information of each frame. To ensure consistent color representation during shape deformation, we propose a continuous color field that constructs a continuous function to represent color values within the geometric space, which is represented by Signed Distance Function (SDF). We decouple shape and color, and design a decoder-only structure to map the shape and color information into two latent vectors. Our coherent 3D shape and color modeling links SDF with color while decoupling color, allowing for independently manipulating color without affecting the shape. Our method does not require an off-the-shelf 2D diffusion model like texture rendering based colored 3D object generation [9]. Next, since direct application of diffusion models to the volumetric SDF and color is computationally intensive, we introduce a matrixized 4D sequence representation, extending the 3D shape and coherent color representations along the temporal dimension. The representations across all frames are jointly concatenated to a matrix capturing the 4D distribution of the sequence. The 4D representation allows the diffusion model to efficiently operate on the temporal dimension, resulting in the generation of intricate 4D sequences with high fidelity. Thirdly, equipped with the efficient 4D representation, we propose an image-text conditioned latent diffusion model for 4D sequence generation. This model captures the intra-frame information, as well as frame-to-frame and frame-to-global relationships, using the proposed hierarchical conditional spatio-temporal attention mechanism to enforce temporal consistency on the content during generation. Our main contributions are:
  • We propose an image-text conditioned 4D volumetric generation framework that enables free navigation with high-quality, consistent rendering.
  • We present a coherent 3D shape and color modeling strategy that ensures consistent color representation during deformation via a continuous color field within geometric space, while enabling flexible shape-color combinations.
  • We introduce a matrixized 4D representation method, independent of pre-rigged pose priors, that facilitates efficient correlation learning along the temporal dimension.
  • We design a hierarchical conditional spatio-temporal attention mechanism to enforce temporal consistency in the generated content during the diffusion process.
We assess unconditioned/conditioned 4D sequences generation across multiple datasets, to show that our approach achieves promising generation performances.
@@ 2 Related Work
3D Generative Models have been extensively explored based on point clouds [2, 27, 46], voxel grids [41, 47, 25], meshes [50], and SDFs [8, 29, 10, 9]. Beyond the generation of 3D shapes, efforts have been made in generating colored objects [9, 28]. Most methods are based on texture rendering, NeRF, score distillation, or Gaussian splatting. Recent works [13, 35, 5, 3, 40] combine the controllability of colored dynamic 3D objects with the expressivity of emerging diffusion models. However, their downsides are long optimization time and poor performance beyond the training views. Arguably, these methods are ineffective for applications involving dynamic signals, e.g. colored 3D moving objects, since the temporal information is not captured.
4D Representations have primarily taken two directions. One approach treats 4D scenes as functions of spatial dimensions (x, y, z) extended by the temporal dimension (t) or latent codes [45, 16, 22]. The alternative approach involves integrating deformation fields with static, canonical 3D models [33, 32, 12, 24]. A key challenge in 4D representation is maintaining computational efficiency along with temporal consistency. Explicit and hybrid representations yield notable improvements in speed and reconstruction quality, e.g., planar decomposition for 4D space-time grids  [6, 15, 38], hash representations [44], and other innovative structures [14, 1] have shown promise. However, current research lacks methods for coherent shape and color representation in 4D spatio-temporal space. While recent advances in volumetric video [19] show promising results, the need for extensive manual stabilization of mesh sequences and the generation of large assets in existing workflows hinders direct volume-based 4D content generation.
4D Generative Models. A prominent research direction employs text-to-video diffusion models to refine 4D representations, exemplified by the optimization of structures like Hexplane [6] or K-plane  [15]. This process typically involves crafting camera trajectories and applying score distillation sampling on the rendered video sequences. Method [3] focuses on enhancing the photorealism of these representations. However, the motion elements are tightly coupled with the 3D content, resulting in lack of diversity and control over the generated motions. An alternate approach uses priors, mostly derived from the skeleton shape parameters and a series of pose parameters to model dynamic sequences [17, 53]. Their motion representations, while plausible, lack compactness and are limited to short time spans. Our method supports the generation of variable-length 4D sequences while allowing for the free combination of shape, motion, and color.
@@ 3 Method
Our objective is to generate volumetric 4D sequence for an object which can provide plausible geometry and appearance at specified time, enabling freely navigate and rendering. Fig. 2 shows an overview of our method. We first introduce a coherent 3D shape and color modeling, which can handle the non-rigid deformation of shape and color details in the 4D sequence. Additionaly, we propose a latent matrixized 4D sequence representation that compresses 4D sequence thus enabling generation. Finnally we introduce a multi-condition diffusion model with proposed Hierarchical Conditional Spatio-Temporal Attention (HCSTA) block to generate the latent matrixized 4D sequence representation under the given conditions. Our method achieves high generation quality while providing exact object at specified times and significantly reduces the occurrence of artifacts compared to other free-form methods, due to its volumetric generation results.
@@ 3.1 Coherent 3D Shape and Color Modeling
The goal of the coherent 3D shape and color modeling is to create an efficient representation of a colored 3D object, and enables the foundation for representing volumetric 4D data. Two elements of the 3D representation are crucial for shape quality and play an important role in the learning of the subsequent 4D stage: (1) the representation should be able to handle the non-rigid deformation of shape and color details in the 4D stage and (2) provide the possibility to learn spatio-temporal coherence at the 4D stage. To achieve these goals, we decompose the shape and color, propossing to use both shape latent representation and color latent representation for coherent 3D shape and color modeling.
@@
During inference, the shape latent code $\hat{\mathbf{z}_{i}}$ for a target SDF $\mathcal{S}^{i}_{gt}$ is estimated via Maximum-a-Posteriori (MAP) as:
@@
Geometric Coherent Color Representation. We then propose a continuous color field within the geometric space to ensure consistent color representation during deformation. Given the predicted $\mathcal{S}^{i}$, we first perform marching cube to reconstruct shape $G_{re}^{i}$ and then use a remesh method [20] to enhance color details by increasing the vertices. As shown in Fig. 3, we align these vertices $v=\{v_{m}\}|^{M}_{m=1}$ and use a KD-tree to extract color values $\mathbf{C}^{i}$ from the ground truth $G_{\mathrm{gt}}^{i}$. These operations leverage the information obtained from the learned shape, indicating that the construction of color representation is constrained by the shape. To ensure that the continuous color field fully represents the color distribution of the object and is generalizable to a broader range of geometries, avoiding fixing the combination of shape and color, we randomly sample points $p_{n}$ both inside and outside the reconstucted surface and use an interpolation function to compute the color values of these points. We construct the continuous color field by: $\mathbf{C}^{i}_{\mathrm{gt}}=\{\Phi_{\theta}(\mathbf{C}^{i},v_{m},p_{n})|n=[1,2,\cdots,N]\}$, where $\Phi_{\theta}$ is the interpolation function. Respectively, we use another latent vector $\mathbf{c}_{i}\in\mathbb{R}^{d_{2}}$ to represente the continuous color field $\mathbf{C}^{i}_{\mathrm{gt}}$ of a 3D object as:
@@
$\mathcal{L}_{c}=\lambda_{1}(L_{1}(\mathbf{C}^{i}_{\mathrm{pred}},$ $\mathbf{C}^{i}_{\mathrm{gt}})+\frac{1}{\sigma^{2}}\|\mathbf{c}_{i}\|_{2}^{2}))$ (5)
$+\lambda_{2}L_{\mathrm{SSIM}}(\mathbf{C}^{i}_{\mathrm{pred}},\mathbf{C}^{i}_{\mathrm{gt}}).$
During inference, the color latent code $\hat{\mathbf{c}_{i}}$ is acquiared using MAP:
$\hat{\mathbf{c}_{i}}=\mathop{\arg\min}_{\mathbf{c}}\displaystyle\sum^{N}_{i=1}\mathcal{L}(C_{\theta}(\mathbf{c}_{i},p_{i}),\mathbf{C}^{i}_{\mathrm{gt}})+\frac{1}{\sigma^{2}}\|\mathbf{c}_{i}\|_{2}^{2}.$ (6)
@@ 3.2 Matrixized Representations in Latent Space
Our goal is to construct a 4D representation optimized for generation. The most straightforward approach is to aggregate $N$ spatial-temporal elements frame-by-frame, leading to time complexity $O(T^{2}N^{2})$ for a sequence with $T$ frames and $N$ elements each. A simplified approach is predicting offsets from a single global code [42, 30, 21] but it loses significant surface geometry, color, and temporal evolution information. The 4D sequence should exhibit smoothness both spatially and temporally, considering the object’s limited elasticity and velocity. Therefore, a latent sets of shape and color is responsible for representing a volumetric 4D sequence. We treat our latent 4D sequence representation as frame stamps on the volumetric 4D sequence, which takes a serial of keyframes for observation. Using our coherent 3D shape and color modeling, we combine multiple 3D shapes into a single 4D representation. We assemble these into a matrix $\mathcal{M}$, capturing 4D dynamics:
@@
$\mathcal{M}=\begin{bmatrix}\mathbf{z}_{1}&\mathbf{z}_{2}&\cdots&\mathbf{z}_{N}\\ \mathbf{c}_{1}&\mathbf{c}_{2}&\cdots&\mathbf{c}_{N}\,\end{bmatrix},$ (7)
where $\mathbf{z}_{i}$, $\mathbf{c}_{i}$ are shape, color latent vectors for frame $i$.
@@
Notably, this latent 4D sequence representation reduces the time complexity to $O(TN^{2})$ in the spatial domain and $O(NT^{2})$ in the temporal domain, while the number of latent elements is significantly smaller than the number of spatial-temporal elements. Compared to single global code methods, $\mathcal{M}$ allows the diffusion model to directly operate on the latent of a specific frame, avoiding drift from single global code and resulting in high-fidelity volumetric 4D sequences. Moreover, this also allows us to apply attention mechanisms to the matrix for shape-to-color supervision, as well as conditional generation. This preserves advantage of our coherent 3D shape and color modeling, namely to allow the shape to supervise the color while also avoiding a fixed combination of shape and color.
Previous works [37, 31] has demonstrated that, when training data is sufficient, latent representations can form a meaningful continuous latent spaces. It’s indicate that although the initial number of frames is fixed, we can use interpolation and MAP estimation to get variable-length 4D sequences:
$\begin{bmatrix}\mathbf{z}_{i}\\ \mathbf{c}_{i}\,\end{bmatrix}=\lambda\begin{bmatrix}\mathbf{z}_{j}\\ \mathbf{c}_{j}\,\end{bmatrix}+(1-\lambda)\begin{bmatrix}\mathbf{z}_{k}\\ \mathbf{c}_{k}\,\end{bmatrix},$ (8)
and MAP estimation:
$\min\mathbb{E}\left[-\sum\log p(\mathbf{z}_{n},\mathbf{c}_{n}\mid\mathbf{z}_{n-k},\mathbf{c}_{n-k})\right].$ (9)
@@
Once the latent representation is generated, our method allows for flexible decoding, enabling the reconstruction of variable-length volumetric 4D sequences by interpolating latent frames in Fig. 2.

### v3

@@ Abstract
keywords:
Generative Model , 4D Generation , Volumetric 4D , 3D Modeling , 4D representation
journal: Pattern Recognition
[…]
[add1]organization=School of Artificial Intelligence, addressline=Xidian University, city=Xi’an, postcode=710126, state=Shanxi, country=China
[…]
[add2]organization=School of Information Engineering, addressline=Chang’an University, city=Xi’an, postcode=710064, state=Shanxi, country=China
[…]
[add3]organization=School of Electrical and Information Engineering, addressline=Hunan University, city=Changsha, postcode=410082, state=Hunan, country=China
[…]
[add4]organization=Computer Science and Software Engineering, addressline=The University of Western Australia, city=Perth, postcode=6009, state=WA, country=Australia
@@ 1 Introduction
Recent works [9, 29] combine the controllability of dynamic 3D objects with the expressivity of emerging diffusion models. 4D representations are proposed to construct an underlying representation of dynamic 3D objects, allowing the diffusion model to learn the distribution of 3D synthesis with an additional temporal dimension. In general, 4D representations can be categorized into prior-based and free-form methods depending on the 3D representation of the output shape. Prior-based methods are mostly derived from shape parameters combined with a series of skeletal pose [8] parameters to model dynamic sequences, e.g., 4D human motion generation [17]. Although they produce plausible results, the generated skeleton human motions assuming a human model prior such as SMPL [22], lacks motion diversity and only supports limited time span. Free-form methods leveraging dynamic Neural Radiance Fields (NeRF) combine the benefits of video and 3D generative models [14], and require heavy computations to generate a 4D NeRF. Moreover, their generated motions are not well-controllable. The 4D Gaussian splatting-based free-form method [15, 29] can produce high-quality renderings, but artifacts often appear when the view changes. This is due to the lack of geometric priors in 4D Gaussian splatting-based free-form approaches. In addition, free-form methods often produce low-quality rendering results during arbitrary navigation. This issue arises because the generated 4D content lacks the essential volumetric properties needed for accurate representation. To this end, we pose a question, how can we efficiently represent a 4D sequence to generate with high-quality geometry, sequence consistency, and alignment with input guidance, while maintaining diversity and allowing for free navigation and rendering?
@@
To address these challenges, we propose several techniques to achieve volumetric 4D sequence generation under given conditions, as shown in Fig. 1. We first design a coherent 3D shape and color modeling method to encode the shape and color information of each frame. To ensure consistent color representation during shape deformation, we propose a continuous color field that constructs a continuous function to represent color values within the geometric space, which is represented by Signed Distance Function (SDF). We decouple shape and color, and design a decoder-only structure to map the shape and color information into two latent vectors. Our coherent 3D shape and color modeling links SDF with color while decoupling color, allowing for independently manipulating color without affecting the shape. Our method does not require an off-the-shelf 2D diffusion model like texture rendering based colored 3D object generation [5]. Next, since direct application of diffusion models to the volumetric SDF and color is computationally intensive, we introduce a matrixized 4D sequence representation, extending the 3D shape and coherent color representations along the temporal dimension. The representations across all frames are jointly concatenated to a matrix capturing the 4D distribution of the sequence. The 4D representation allows the diffusion model to efficiently operate on the temporal dimension, resulting in the generation of intricate 4D sequences with high fidelity. Thirdly, equipped with the efficient 4D representation, we propose an image-text conditioned latent diffusion model for 4D sequence generation. This model captures the intra-frame information, as well as frame-to-frame and frame-to-global relationships, using the proposed hierarchical conditional spatio-temporal attention mechanism to enforce temporal consistency on the content during generation.
Our main contributions are: (1) We propose an image-text conditioned 4D volumetric generation framework that enables free navigation with high-quality, consistent rendering. (2) We present a coherent 3D shape and color modeling strategy that ensures consistent color representation during deformation via a continuous color field within geometric space, while enabling flexible shape-color combinations. (3) We introduce a matrixized 4D representation method, independent of pre-rigged pose priors, that facilitates efficient correlation learning along the temporal dimension. (4) We design a hierarchical conditional spatio-temporal attention mechanism to enforce temporal consistency in the generated content during the diffusion process. We assess unconditioned/conditioned 4D sequences generation across multiple datasets, to show that our approach achieves promising generation performances.
@@ 2 Related Work
3D Generative Models have been extensively explored based on point clouds [37], voxel grids [21], meshes [40], and SDFs [5]. Beyond the generation of 3D shapes, efforts have been made in generating colored objects [5, 23]. Most methods are based on texture rendering, NeRF, score distillation, or Gaussian splatting. Recent works [9, 29] combine the controllability of colored dynamic 3D objects with the expressivity of emerging diffusion models. However, their downsides are long optimization time and poor performance beyond the training views. Arguably, these methods are ineffective for applications involving dynamic signals, e.g. colored 3D moving objects, since the temporal information is not captured.
4D Representations have primarily taken two directions. One approach treats 4D scenes as functions of spatial dimensions (x, y, z) extended by the temporal dimension (t) or latent codes [36, 19]. The alternative approach involves integrating deformation fields with static, canonical 3D models [27, 24]. A key challenge in 4D representation is maintaining computational efficiency along with temporal consistency. Explicit and hybrid representations yield notable improvements in speed and reconstruction quality, e.g., planar decomposition for 4D space-time grids  [11], hash representations [35], and other innovative structures [1] have shown promise. However, current research lacks methods for coherent shape and color representation in 4D spatio-temporal space. While recent advances in volumetric video [13] show promising results, the need for extensive manual stabilization of mesh sequences and the generation of large assets in existing workflows hinders direct volume-based 4D content generation.
4D Generative Models. A prominent research direction employs text-to-video diffusion models to refine 4D representations, exemplified by the optimization of structures like Hexplane [3] or K-plane  [10]. This process typically involves crafting camera trajectories and applying score distillation sampling on the rendered video sequences. Method [2] focuses on enhancing the photorealism of these representations. However, the motion elements are tightly coupled with the 3D content, resulting in lack of diversity and control over the generated motions. An alternate approach uses priors, mostly derived from the skeleton shape parameters and a series of pose parameters to model dynamic sequences [8]. Their motion representations, while plausible, lack compactness and are limited to short time spans. Our method supports the generation of variable-length 4D sequences while allowing for the free combination of shape, motion, and color.
@@ 3 Method
Our objective is to generate volumetric 4D sequence for an object which can provide plausible geometry and appearance at specified time, enabling freely navigate and rendering. Fig. 2 shows an overview of our method. We first introduce a coherent 3D shape and color modeling, which can handle the non-rigid deformation of shape and color details in the 4D sequence. Additionally, we propose a latent matrixized 4D sequence representation that compresses 4D sequence thus enabling generation. Finally we introduce a multi-condition diffusion model with proposed Hierarchical Conditional Spatio-Temporal Attention (HCSTA) block to generate the latent matrixized 4D sequence representation under the given conditions. Our method achieves high generation quality while providing exact object at specified times and significantly reduces the occurrence of artifacts compared to other free-form methods, due to its volumetric generation results.
@@ 3.1 Coherent 3D Shape and Color Modeling
The goal of the coherent 3D shape and color modeling is to create an efficient representation of a colored 3D object, and enables the foundation for representing volumetric 4D data. Two elements of the 3D representation are crucial for shape quality and play an important role in the learning of the subsequent 4D stage: (1) the representation should be able to handle the non-rigid deformation of shape and color details in the 4D stage and (2) provide the possibility to learn spatio-temporal coherence at the 4D stage. To achieve these goals, we decompose the shape and color, proposing to use both shape latent representation and color latent representation for coherent 3D shape and color modeling, as shown in Fig. 3.
@@
During inference, the shape latent code $\hat{\mathbf{z}_{i}}$ for a target SDF $\mathcal{S}^{i}_{gt}$ is estimated via Maximum-a-Posterior (MAP) as:
@@
Geometric Coherent Color Representation. We then propose a continuous color field within the geometric space to ensure consistent color representation during deformation. Given the predicted $\mathcal{S}^{i}$, we first perform marching cube to reconstruct shape $G_{re}^{i}$ and then use a remesh method [16] to enhance color details by increasing the vertices. As shown in Fig. 3, we align these vertices $v=\{v_{m}\}|^{M}_{m=1}$ and use a KD-tree to extract color values $\mathbf{C}^{i}$ from the ground truth $G_{\mathrm{gt}}^{i}$. These operations leverage the information obtained from the learned shape, indicating that the construction of color representation is constrained by the shape. To ensure that the continuous color field fully represents the color distribution of the object and is generalizable to a broader range of geometries, avoiding fixing the combination of shape and color, we randomly sample points $p_{n}$ both inside and outside the reconstructed surface and use an interpolation function to compute the color values of these points. We construct the continuous color field by: $\mathbf{C}^{i}_{\mathrm{gt}}=\{\Phi_{\theta}(\mathbf{C}^{i},v_{m},p_{n})|n=[1,2,\cdots,N]\}$, where $\Phi_{\theta}$ is the interpolation function. Respectively, we use another latent vector $\mathbf{c}_{i}\in\mathbb{R}^{d_{2}}$ to represent the continuous color field $\mathbf{C}^{i}_{\mathrm{gt}}$ of a 3D object as:
@@
$\mathcal{L}_{c}=\lambda_{1}(L_{1}(\mathbf{C}^{i}_{\mathrm{pred}},\mathbf{C}^{i}_{\mathrm{gt}})+\frac{1}{\sigma^{2}}\|\mathbf{c}_{i}\|_{2}^{2}))+\lambda_{2}L_{\mathrm{SSIM}}(\mathbf{C}^{i}_{\mathrm{pred}},\mathbf{C}^{i}_{\mathrm{gt}}).$ (5)
During inference, the color latent code $\hat{\mathbf{c}_{i}}$ is acquired using MAP:
$\hat{\mathbf{c}_{i}}=\mathop{\arg\min}_{\mathbf{c}}\displaystyle\sum^{N}_{i=1}\mathcal{L}_{c}(C_{\theta}(\mathbf{c}_{i},p_{i}),\mathbf{C}^{i}_{\mathrm{gt}})+\frac{1}{\sigma^{2}}\|\mathbf{c}_{i}\|_{2}^{2}.$ (6)
@@ 3.2 Matrixized Representations in Latent Space
Our goal is to construct a 4D representation optimized for generation. The most straightforward approach is to aggregate $N$ spatial-temporal elements frame-by-frame, leading to time complexity $O(T^{2}N^{2})$ for a sequence with $T$ frames and $N$ elements each, which is computationally infeasible. A simplified approach is predicting offsets from a single global code [34, 25, 18] but it loses significant surface geometry, color, and temporal evolution information. The 4D sequence should exhibit smoothness both spatially and temporally, considering the limited elasticity and velocity of object. Therefore, a continuous latent sets of shape and color is responsible for representing a volumetric 4D sequence. Building upon our Coherent 3D Shape and Color Modeling, our latent representations can form a meaningful continuous latent space when training data is sufficient [31, 26], exhibiting spatiotemporal continuity. Based on spatiotemporal continuity, we propose a keyframe-based Matrixized 4D Representation.
We treat our latent 4D sequence representation as frame stamps on the volumetric 4D sequence, which takes a serial of keyframes for observation. Based on our interpolatable and meaningful continuous latent space, we propose a $K$-controlled matrixized 4D representation:
@@
$\mathcal{M}=\begin{bmatrix}\mathbf{z}_{1}&\overbrace{[\mathbf{z}_{u1\xrightarrow{}\cdots}]}^{K\text{ frames}}&\mathbf{z}_{2}&\overbrace{[\mathbf{z}_{u2\xrightarrow{}\cdots}]}^{K\text{ frames}}&\cdots&\overbrace{[\mathbf{z}_{un-1\xrightarrow{}\cdots}]}^{K\text{ frames}}&\mathbf{z}_{T}\\ \mathbf{c}_{1}&\underbrace{[\mathbf{c}_{u1\xrightarrow{}\cdots}]}_{K\text{ frames}}&\mathbf{c}_{2}&\underbrace{[\mathbf{c}_{u2\xrightarrow{}\cdots}]}_{K\text{ frames}}&\cdots&\underbrace{[\mathbf{c}_{un-1\xrightarrow{}\cdots}]}_{K\text{ frames}}&\mathbf{c}_{T}\,\end{bmatrix},$ (7)
where $\mathbf{z}_{i}$, $\mathbf{c}_{i}$ are shape and color latent vectors for keyframe $i$, $[\mathbf{z}_{ui\rightarrow\cdots}]$, $[\mathbf{c}_{ui\rightarrow\cdots}]$ denote the unseen $K$ frames in the proposed matrixized representation, which can be recovered via interpolation from adjacent frames to balance efficiency and accuracy. We determine $K$ by evaluating the quality of unseen frames using the reconstruction loss defined in our coherent modeling on the interpolated 4D sequences:
$\arg$ $\max\limits_{K}\mathbb{E}\left[\mathcal{L}_{s}(\psi(\mathbf{z}_{1}\cdots\mathbf{z}_{T}))+\mathcal{L}_{c}(\psi(\mathbf{c}_{1}\cdots\mathbf{c}_{T}))-\mathcal{L}_{s}(\hat{\mathbf{z}}_{1}\cdots\hat{\mathbf{z}}_{T})-\mathcal{L}_{c}(\hat{\mathbf{c}}_{1}\cdots\hat{\mathbf{c}}_{T})\right].$ (8)
[…]
Here, $(\hat{\mathbf{z}}_{1}\cdots\hat{\mathbf{z}}_{T})$ and $(\hat{\mathbf{c}}_{1}\cdots\hat{\mathbf{c}}_{T})$ are the original latents of 4D sequence, $\sigma$ is a small value computed as a proportion of $\mathbb{E}\left[\mathcal{L}_{s}(\hat{\mathbf{z}}_{1},\cdots,\hat{\mathbf{z}}_{T})+\mathcal{L}_{c}(\hat{\mathbf{c}}_{1},\cdots,\hat{\mathbf{c}}_{T})\right]$, $\psi$ denotes an interpolation function for recovering $K$ unseen frames, which can be implemented using linear interpolation
$\begin{bmatrix}[\mathbf{z}_{ui\xrightarrow{}\cdots}]\\ [\mathbf{c}_{ui\xrightarrow{}\cdots}]\end{bmatrix}=\Lambda^{T}\begin{bmatrix}\mathbf{z}_{i}\\ \mathbf{c}_{i}\,\end{bmatrix}+(1-\Lambda^{T})\begin{bmatrix}\mathbf{z}_{i+1}\\ \mathbf{c}_{i+1}\,\end{bmatrix},$ (10)
through maximum a posterior (MAP) estimation:
$\min\mathbb{E}\left[-\sum\log p([\mathbf{z}_{ui\xrightarrow{}\cdots}],[\mathbf{c}_{ui\xrightarrow{}\cdots}]\mid\mathbf{z}_{i},\mathbf{c}_{i},\mathbf{z}_{i+1},\mathbf{c}_{i+1})\right].$ (11)
Notably, this latent 4D sequence representation reduces the complexity to $O(\frac{T^{2}}{K^{2}}d)$, where $d$ is the dimension of $\mathbf{z}$ and $\mathbf{c}$. Typically, for a $5$-second sequence, $K$ ranges from $10$ to $20$, while $T$ is around 120. Compared to methods that assemble every frame individually with complexity $O(T^{2}d)$, our approach uses $\frac{T^{2}}{K^{2}}\ll T^{2}$ and $d\ll N$, further demonstrating the efficiency of our 4D representation.
@@
Compared to single global code methods, $\mathcal{M}$ allows the diffusion model to directly operate on the latent of a specific frame, avoiding drift from single global code and resulting in high-fidelity volumetric 4D sequences. Moreover, this also allows us to apply attention mechanisms to the matrix for shape-to-color supervision, as well as conditional generation. This preserves advantage of our coherent 3D shape and color modeling, namely to allow the shape to supervise the color while also avoiding a fixed combination of shape and color.
@@
Our matrixized 4D representations indicate that, although the initial number of keyframes for one 4D sequence is fixed, variable-length 4D sequences can be obtained from two or more 4D sequences through recovering the unseen frames via Eq. 9 and  10. Once the latent representation is generated, our method allows for flexible decoding, enabling the reconstruction of variable-length volumetric 4D sequences by interpolating latent keyframes in Fig. 2.